# Onboard Mission Replanning for Adaptive Cooperative Multi-Robot Systems

### v3

@@ Abstract
©2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works. DOI: TBC
@@ I Introduction
Teams of autonomous robots, including uncrewed aerial, underwater, surface, or ground vehicles (UAVs, UUVs, USVs, UGVs) and multi-satellite constellations, have huge potential for efficiently and safely completing a diverse range of complex missions. In comparison to single-agent operations, such multi-agent systems have increased robustness to failure [1], greater areal coverage [2], and faster mission completion [3]. Furthermore, they can cooperatively tackle complex missions comprised of multiple tasks, where these tasks can be as diverse as capturing images [4], manufacturing/ repairing items or transporting materials [5], or making environmental measurements [6]. However, these systems also commonly operate in environments that are both remote or isolated and also highly dynamic and hazardous, e.g. in nuclear power plants [7], over active volcanoes [6] or in congested low earth orbits in space [4]. Operating in such challenging environments requires these systems to be able to adapt and replan in response to major unexpected changes that impact the mission plan [8] such as the identification of new tasks or the loss or failure of a robotic agent, without relying on fragile or slow communication links to centralized ground stations [9] (Fig. 1a). Therefore, lightweight and fast replanning algorithms that can be deployed on-board cooperative robotic systems are critical for enabling autonomous adaptability during a mission; but these do not yet exist.
@@
[10] TSP RNN Ptr. Net. REINFORCE $\times$ $\times$ $\times$ $\times$ $\times$
[11] TSP Attention Ptr. Net. REINFORCE (Rollout) $\times$ $\times$ $\times$ $\times$ $\times$
[12] TSP RNN + Graph Attention Ptr. Net. REINFORCE (Rollout) $\times$ $\times$ $\times$ $\times$ $\times$
[13] CVRP Graph Attention Attention Model REINFORCE (Rollout) $\times$ $\times$ $\times$ $\times$ $\times$
[14] multi-tasks VRP Graph Attention Attention Model POMO + MoE Layer $\times$ $\times$ $\times$ ✓(Goal) $\times$
[15] mTSP with Multiple Depots Graph Attention Ptr. Net. REINFORCE $\times$ $\times$ $\times$ $\times$
[16] mTSP with Multiple Depots MoE Net. Ptr. Net. Pretrained $\times$ $\times$ ✓(Agent)
[17] Split Delivery VRP Convolutional Ptr. Net. REINFORCE $\times$ $\times$ $\times$
[18] VRP (Min-Max) Graph Attention (Partition/Navigation) Attention Model REINFORCE (Rollout) $\times$ $\times$ ✓(Depot) $\times$
[19] Heterogeneous CVRP (Min-Max) Graph Attention Attention Model (Parallel) REINFORCE $\times$ $\times$ ✓(Agent) $\times$
@@
Because mission planning with multiple agents requires finding the optimal tour path for each agent while minimizing the length of the longest tour, previous studies have often formulated and solved it as a variant of the multiple Traveling Salesman Problem (mTSP) with a min-max objective [20]. However, our particular planning problem has three major differences to traditional mTSP (Fig. 1b), and in addition requires practical consideration for edge deployment onboard robotic systems. First, variable start locations for agents must be considered, as each agent may be at any location within the mission domain when replanning is required. Second, nodes with variable time costs are necessary to reflect the common real-world situation where mission time depends not just on how long it takes for an agent to travel to a task, but also how long it takes to complete that task. This models a wide range of applications where tasks have differing execution times [21, 22, 23]. Third, cooperative tasking is required, such that multiple agents can work together on a single task to complete it faster, even if this joint contribution is asynchronous rather than simultaneous. No previous studies have addressed this important combination of multiple adaptations (see “Key Adaptation” columns of Table I), which we name the Cooperative Mission Replanning Problem (CMRP). Finally, to enable onboard replanning, any model for solving CMRP must be suitably small and fast [24]. These limitations also necessitate a single generalized model that can solve a range of problem sizes, as the number of available agents and remaining tasks are likely to vary during a mission [25] and it is impractical to store one model per problem size on edge devices.
@@
The mTSP class of problems are NP-hard [26] with the total number of possible mission plans scaling as $\frac{\left(n+m-1\right)!}{2\left(m-1\right)!}$, where $n$ is the number of tasks and $m$ is the number of agents. Reinforcement Learning (RL) techniques, especially those using Attention mechanisms (e.g. GAT; Graph Attention Networks), have proven to be effective at overcoming these scaling issues and efficiently solving TSP [10, 11, 12], mTSP [27], min-max Vehicle Routing Problems (VRP) [18, 19], or other similar multi-agent problems [13, 14].
@@
These methods are able to scale to larger problems than exact solvers such as CPLEX [28], are faster at execution time and generalize better than heuristic methods [29, 27, 30] such as the Lin-Kernighan heuristic [31], show more consistent and generalizable performance than metaheuristic methods [32, 33], and can be deployed on edge devices as lightweight models [34].
@@
But despite its real-world importance, no work to date has addressed the replanning problem as outlined above. Several studies have made one of the required adaptations, e.g. multiple depots, where agents start and finish at different locations [15, 16, 18] or cooperative tasking [17], but none have considered more than one together, and no previous studies have considered nodes with variable time costs (Table I), which is a fundamental requirement. Similarly, most prior research has focused on high-performance computing systems with GPUs, with limited investigation of deployment on ground robots [16] or CPUs [17]. And whilst [14, 17, 18, 19], developed a single neural solver for different problem variants, numbers of agents or depots, they did not consider varying numbers of tasks. Our work fills this critical gap, developing a cooperative replanning solver that is suitable for edge deployment and real-world on-board application.
@@ II-A The Cooperative Mission Replanning Problem (CMRP)
Traditional mTSP with $m$ salespeople or agents and $n$ cities or tasks can be defined as a graph $G=(V,E)$ where the vertices $V=\{v_{depot},v_{1},...v_{n}\}$ include all $n$ task locations as well as $v_{depot}$, which serves as the single start and end location for all agents. The edge set $E$ is defined as $E=\{(v_{i},v_{j}):v_{i},v_{j}\in V,i\neq j\}$ and represents all possible valid paths between tasks. For all edges, $T=(d_{ij})$ represents the travel time associated with traversing from task $v_{i}$ to $v_{j}$, assuming a constant travel speed of 11 m/s.
First, we incorporate flexible start locations into this formulation by adding $m$ start locations to the vertex set, similar to the multiple depot problems [15, 16]. Our new vertex set is $V^{\prime}=\{v_{depot},v_{start1},...v_{startm},v_{1},...v_{n}\}$, where $v_{startk}$ is the start location of the $k$-th agent, and $v_{i}$ is the location of the $i$-th task. Moreover, unlike formulations in the literature where agents complete their tours at their respective starting locations, our $v_{depot}$ serves as a common destination for all agents, simulating a shared landing space for drones. Two constraints are set to ensure that agents always begin at their designated start locations: (1) $d_{depot\ i}=0$; (2) $d_{i\ startj}=\infty$; where $v_{i}\in\{v_{1},...v_{n}\}$ and $v_{startj}\in\{v_{start1},...v_{startm}\}$. Hence, the $T$ cost matrix is no longer symmetric and therefore changes the problem to an asymmetric mTSP.
Second, we incorporate variable time costs for each task so that each vertex $v_{i}$ has an associated non-negative time cost $t_{i}$, with the exception of the depot and the start locations, where $t_{i}=0\text{ for }v_{i}\in\{v_{depot},v_{start1},...v_{startm}\}$. This is similar to the nodes’ demand in Capacitated VRP (CVRP) [13] with the distinction that $t_{i}$ is incorporated into the edge costs rather than being a separate vector. Our new cost matrix is therefore calculated as $T^{\prime}=d_{ij}+t_{j}$, and again this modification introduces more asymmetry to this matrix, as time costs are no longer determined just by the distance between two vertices, but also by the direction of travel.
Third, each task is discretized into $\delta$ sub-tasks similar to the work in Split Delivery VRP [17] to enable cooperative tasking, such that the total number of sub-tasks $n^{\prime}=n\times\delta$. We refer to $\delta$ as the task discretization level. Sub-tasks replace tasks in the vertex set. They have the same spatial position as the original task but contain a corresponding fraction of the original task completion time. The time cost for a sub-task associated with task $i$ is $t^{\prime}_{i}=t_{i}/\delta$.
@@ II-B The Graph Attention Replanner (GATR)
Our proposed GAT-based mission replanner (GATR) builds upon the encoder-decoder architecture from [13]. We adapt the mTSP environment from the RL4CO Library [27] to incorporate different agents’ start locations and time costs associated with nodes (the first and second changes to standard mTSP described in Table I). Discretization of tasks to allow splitting of tasks between cooperative agents (the third change) does not require any additional adaptation, and can simply be accommodated during problem setup by replacing a single task of a certain time cost with several shorter collocated tasks. Finally, we enhance the data ingestion pipeline to enable training of a single generalized model for multiple problem sizes.
The overall training process is shown in Fig. 2a. Using the environment described above, we trained our model using the REINFORCE algorithm with a greedy rollout baseline [13] (see Appendix). The reward is defined as the negative of the maximum (largest) mission time among the agents as in [18].
@@ II-B1 Encoder
The primary role of the GAT encoder is to transform an input graph into latent representations, as illustrated in the first step of the flowchart in Fig. 2a. In addition to the standard mTSP graph features (depot locations and task locations), we also include the time cost of vertices and agents’ starting locations as part of the input to the encoder. This is similar to the approach used for specifying the ‘demand’ for deliveries in the CVRP [13], and for incorporating multiple depots in [15]. The GAT encoder then transforms these features into a set of node embeddings and a graph embedding that are passed to the decoder using the procedure described in the Appendix.
@@ II-B2 Decoder
Our decoder is based on the Attention Model, which was selected over the traditional Pointer Network due to its empirically superior performance [12]. The decoder forms the mission plan sequentially, as illustrated in Fig. 2c-d. At each time step, the decoder selects the next sub-task to visit for an agent based on the graph embedding, the current node embedding, the current state of the mission (in embedding space) and the availability mask. If the depot is selected, the algorithm proceeds to the subsequent agent. Note that an agent has the option to stop completing tasks and return to the depot at any point, provided there is still at least one other agent available to complete remaining tasks, similar to the assumption made in [16]. The state of the environment is summarized by the following features: the number of remaining agents, the accumulated mission time for the current agent, the maximum mission time across all agents, and the distance from the depot. These features are mapped into an embedding space using affine linear transformations and concatenated with the embedding for the current agent vertex (i.e. the agent’s current location) to form the context embeddings. Unlike in [13], we do not include the embedding of the first vertex traversed by the agent, as the agent does not return to this location in the replanning problem. Cross-attention with an availability mask is applied to the context embeddings and vertex embeddings to compute the attention scores. In the replanning problem, the mask prevents selection of: previously visited sub-tasks; start locations when the current node is not the depot; and the depot if the current agent is the last one available, and unvisited sub-tasks remain. The next sub-task is then selected based on these scores, as detailed in the Appendix. This process is repeated until all sub-tasks are assigned to the agents. The final output is a set of routes, one for each agent, covering all sub-tasks, as illustrated in Fig. 2d. We note that in future work, the GATR architecture could be easily extended, e.g. to other encoders tailored for min-max objectives (e.g. [18], which has separate embeddings for task partitioning and navigation), or decoders that output actions for agents in parallel to encourage collaborative behavior [19].
@@ III-A Analysis for small problem sizes
We evaluate mission times against four benchmarks: the optimal and median population solutions (Opt-P and Med-P respectively), calculated via brute-force enumeration of the full population of possible solutions (e.g., 1,814,400 for 3 agents and 8 sub-tasks, from Equation 1, evaluation of which takes $\sim$1 minute for a single scenario on the GPU); the median of a sample of the population - 101 randomly-generated solutions (Med-S); and a near-optimal solution generated by the LKH3 solver [31] (NOpt-LKH3). Opt-P and NOpt-LKH3 serve as the upper bounds on expected performance, whilst the two median benchmarks represent a lower bound as they simulate the expected performance of a randomly generated plan. We note that although it is feasible to obtain enumeration solutions for these small problem sizes, this is not possible for the larger problems we investigate in Section III-B, for which NOpt-LKH3 and Med-S are essential for evaluation instead.

### v4

@@ Abstract
©2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works. Link to published paper: https://ieeexplore.ieee.org/document/11222773. DOI: 10.1109/LRA.2025.3627093
@@ I Introduction
Teams of autonomous robots, including uncrewed aerial, underwater, surface, or ground vehicles (UAVs, UUVs, USVs, UGVs) and multi-satellite constellations, have huge potential for efficiently and safely completing a diverse range of complex missions. In comparison to single-agent operations, such multi-agent systems have increased robustness to failure [safety_in_number], greater areal coverage [coverage], and faster mission completion [thakur_swarm_2021]. Furthermore, they can cooperatively tackle complex missions comprised of multiple tasks, where these tasks can be as diverse as capturing images [satellite1], manufacturing/ repairing items or transporting materials [node_cost], or making environmental measurements [palma]. However, these systems also commonly operate in environments that are both remote or isolated and also highly dynamic and hazardous, e.g. in nuclear power plants [powerengineering], over active volcanoes [palma] or in congested low earth orbits in space [satellite1]. Operating in such challenging environments requires these systems to be able to adapt and replan in response to major unexpected changes that impact the mission plan [rizk_cooperative_2020] such as the identification of new tasks or the loss or failure of a robotic agent, without relying on fragile or slow communication links to centralized ground stations [live-fly] (Fig. 1a). Therefore, lightweight and fast replanning algorithms that can be deployed on-board cooperative robotic systems are critical for enabling autonomous adaptability during a mission; but these do not yet exist.
@@
[bello2017neuralcombinatorialoptimizationreinforcement] TSP RNN Ptr. Net. REINFORCE $\times$ $\times$ $\times$ $\times$ $\times$
[deudon] TSP Attention Ptr. Net. REINFORCE (Rollout) $\times$ $\times$ $\times$ $\times$ $\times$
[hgpn] TSP RNN + Graph Attention Ptr. Net. REINFORCE (Rollout) $\times$ $\times$ $\times$ $\times$ $\times$
[kool2019attentionlearnsolverouting] CVRP Graph Attention Attention Model REINFORCE (Rollout) $\times$ $\times$ $\times$ $\times$ $\times$
[mvmoe] multi-tasks VRP Graph Attention Attention Model POMO + MoE Layer $\times$ $\times$ $\times$ ✓(Goal) $\times$
[shi_multi-agent_2023] mTSP with Multiple Depots Graph Attention Ptr. Net. REINFORCE $\times$ $\times$ $\times$ $\times$
[ucl_multi_robot] mTSP with Multiple Depots MoE Net. Ptr. Net. Pretrained $\times$ $\times$ ✓(Agent)
[nazari2018reinforcementlearningsolvingvehicle] Split Delivery VRP Convolutional Ptr. Net. REINFORCE $\times$ $\times$ $\times$
[dpn] VRP (Min-Max) Graph Attention (Partition/Navigation) Attention Model REINFORCE (Rollout) $\times$ $\times$ ✓(Depot) $\times$
[parco] Heterogeneous CVRP (Min-Max) Graph Attention Attention Model (Parallel) REINFORCE $\times$ $\times$ ✓(Agent) $\times$
@@
Because mission planning with multiple agents requires finding the optimal tour path for each agent while minimizing the length of the longest tour, previous studies have often formulated and solved it as a variant of the multiple Traveling Salesman Problem (mTSP) with a min-max objective [mtsp]. However, our particular planning problem has three major differences to traditional mTSP (Fig. 1b), and in addition requires practical consideration for edge deployment onboard robotic systems. First, variable start locations for agents must be considered, as each agent may be at any location within the mission domain when replanning is required. Second, nodes with variable time costs are necessary to reflect the common real-world situation where mission time depends not just on how long it takes for an agent to travel to a task, but also how long it takes to complete that task. This models a wide range of applications where tasks have differing execution times [aggarwal_extended_2022, uav_remote_sensing, uav_data_collection]. Third, cooperative tasking is required, such that multiple agents can work together on a single task to complete it faster, even if this joint contribution is asynchronous rather than simultaneous. No previous studies have addressed this important combination of multiple adaptations (see “Key Adaptation” columns of Table I), which we name the Cooperative Mission Replanning Problem (CMRP). Finally, to enable onboard replanning, any model for solving CMRP must be suitably small and fast [elim_bnn]. These limitations also necessitate a single generalized model that can solve a range of problem sizes, as the number of available agents and remaining tasks are likely to vary during a mission [changing_number] and it is impractical to store one model per problem size on edge devices.
@@
The mTSP class of problems are NP-hard [nphard] with the total number of possible mission plans scaling as $\frac{\left(n+m-1\right)!}{2\left(m-1\right)!}$, where $n$ is the number of tasks and $m$ is the number of agents. Reinforcement Learning (RL) techniques, especially those using Attention mechanisms (e.g. GAT; Graph Attention Networks), have proven to be effective at overcoming these scaling issues and efficiently solving TSP [bello2017neuralcombinatorialoptimizationreinforcement, deudon, hgpn], mTSP [rl4co], min-max Vehicle Routing Problems (VRP) [dpn, parco], or other similar multi-agent problems [kool2019attentionlearnsolverouting, mvmoe].
@@
These methods are able to scale to larger problems than exact solvers such as CPLEX [cplex2009v12], are faster at execution time and generalize better than heuristic methods [surveyrl, rl4co, surveyml] such as the Lin-Kernighan heuristic [lkh3], show more consistent and generalizable performance than metaheuristic methods [turner_fast_2018, Li_2020], and can be deployed on edge devices as lightweight models [8947146].
@@
But despite its real-world importance, no work to date has addressed the replanning problem as outlined above. Several studies have made one of the required adaptations, e.g. multiple depots, where agents start and finish at different locations [shi_multi-agent_2023, ucl_multi_robot, dpn] or cooperative tasking [nazari2018reinforcementlearningsolvingvehicle], but none have considered more than one together, and no previous studies have considered nodes with variable time costs (Table I), which is a fundamental requirement. Similarly, most prior research has focused on high-performance computing systems with GPUs, with limited investigation of deployment on ground robots [ucl_multi_robot] or CPUs [nazari2018reinforcementlearningsolvingvehicle]. And whilst [mvmoe, nazari2018reinforcementlearningsolvingvehicle, dpn, parco], developed a single neural solver for different problem variants, numbers of agents or depots, they did not consider varying numbers of tasks. Our work fills this critical gap, developing a cooperative replanning solver that is suitable for edge deployment and real-world on-board application.
@@ II-A The Cooperative Mission Replanning Problem (CMRP)
Traditional mTSP with $m$ salespeople or agents and $n$ cities or tasks can be defined as a graph $G=(V,E)$ where the vertices $V=\{v_{depot},v_{1},...v_{n}\}$ include all $n$ task locations as well as $v_{depot}$, which serves as the single start and end location for all agents. The edge set $E$ is defined as $E=\{(v_{i},v_{j}):v_{i},v_{j}\in V,i\neq j\}$ and represents all possible valid paths between tasks. For all edges, $T=(d_{ij})$ represents the travel time associated with traversing from task $v_{i}$ to $v_{j}$, assuming a constant travel speed of $1$ m/s.
First, we incorporate flexible start locations into this formulation by adding $m$ start locations to the vertex set, similar to the multiple depot problems [shi_multi-agent_2023, ucl_multi_robot]. Our new vertex set is $V^{\prime}=\{v_{depot},v_{start1},...v_{startm},v_{1},...v_{n}\}$, where $v_{startk}$ is the start location of the $k$-th agent, and $v_{i}$ is the location of the $i$-th task. Moreover, unlike formulations in the literature where agents complete their tours at their respective starting locations, our $v_{depot}$ serves as a common destination for all agents, simulating a shared landing space for drones. Two constraints are set to ensure that agents always begin at their designated start locations: (1) $d_{depot\ i}=0$; (2) $d_{i\ startj}=\infty$; where $v_{i}\in\{v_{1},...v_{n}\}$ and $v_{startj}\in\{v_{start1},...v_{startm}\}$. Hence, the $T$ cost matrix is no longer symmetric and therefore changes the problem to an asymmetric mTSP.
Second, we incorporate variable time costs for each task so that each vertex $v_{i}$ has an associated non-negative time cost $t_{i}$, with the exception of the depot and the start locations, where $t_{i}=0\text{ for }v_{i}\in\{v_{depot},v_{start1},...v_{startm}\}$. This is similar to the nodes’ demand in Capacitated VRP (CVRP) [kool2019attentionlearnsolverouting] with the distinction that $t_{i}$ is incorporated into the edge costs rather than being a separate vector. Our new cost matrix is therefore calculated as $T^{\prime}=d_{ij}+t_{j}$, and again this modification introduces more asymmetry to this matrix, as time costs are no longer determined just by the distance between two vertices, but also by the direction of travel.
Third, each task is discretized into $\delta$ sub-tasks similar to the work in Split Delivery VRP [nazari2018reinforcementlearningsolvingvehicle] to enable cooperative tasking, such that the total number of sub-tasks $n^{\prime}=n\times\delta$. We refer to $\delta$ as the task discretization level. Sub-tasks replace tasks in the vertex set. They have the same spatial position as the original task but contain a corresponding fraction of the original task completion time. The time cost for a sub-task associated with task $i$ is $t^{\prime}_{i}=t_{i}/\delta$.
@@ II-B The Graph Attention Replanner (GATR)
Our proposed GAT-based mission replanner (GATR) builds upon the encoder-decoder architecture from [kool2019attentionlearnsolverouting]. We adapt the mTSP environment from the RL4CO Library [rl4co] to incorporate different agents’ start locations and time costs associated with nodes (the first and second changes to standard mTSP described in Table I). Discretization of tasks to allow splitting of tasks between cooperative agents (the third change) does not require any additional adaptation, and can simply be accommodated during problem setup by replacing a single task of a certain time cost with several shorter collocated tasks. Finally, we enhance the data ingestion pipeline to enable training of a single generalized model for multiple problem sizes.
The overall training process is shown in Fig. 2a. Using the environment described above, we trained our model using the REINFORCE algorithm with a greedy rollout baseline [kool2019attentionlearnsolverouting] (see Appendix). The reward is defined as the negative of the maximum (largest) mission time among the agents as in [dpn].
@@ II-B1 Encoder
The primary role of the GAT encoder is to transform an input graph into latent representations, as illustrated in the first step of the flowchart in Fig. 2a. In addition to the standard mTSP graph features (depot locations and task locations), we also include the time cost of vertices and agents’ starting locations as part of the input to the encoder. This is similar to the approach used for specifying the ‘demand’ for deliveries in the CVRP [kool2019attentionlearnsolverouting], and for incorporating multiple depots in [shi_multi-agent_2023]. The GAT encoder then transforms these features into a set of node embeddings and a graph embedding that are passed to the decoder using the procedure described in the Appendix.
@@ II-B2 Decoder
Our decoder is based on the Attention Model, which was selected over the traditional Pointer Network due to its empirically superior performance [hgpn]. The decoder forms the mission plan sequentially, as illustrated in Fig. 2c-d. At each time step, the decoder selects the next sub-task to visit for an agent based on the graph embedding, the current node embedding, the current state of the mission (in embedding space) and the availability mask. If the depot is selected, the algorithm proceeds to the subsequent agent. Note that an agent has the option to stop completing tasks and return to the depot at any point, provided there is still at least one other agent available to complete remaining tasks, similar to the assumption made in [ucl_multi_robot]. The state of the environment is summarized by the following features: the number of remaining agents, the accumulated mission time for the current agent, the maximum mission time across all agents, and the distance from the depot. These features are mapped into an embedding space using affine linear transformations and concatenated with the embedding for the current agent vertex (i.e. the agent’s current location) to form the context embeddings. Unlike in [kool2019attentionlearnsolverouting], we do not include the embedding of the first vertex traversed by the agent, as the agent does not return to this location in the replanning problem. Cross-attention with an availability mask is applied to the context embeddings and vertex embeddings to compute the attention scores. In the replanning problem, the mask prevents selection of: previously visited sub-tasks; start locations when the current node is not the depot; and the depot if the current agent is the last one available, and unvisited sub-tasks remain. The next sub-task is then selected based on these scores, as detailed in the Appendix. This process is repeated until all sub-tasks are assigned to the agents. The final output is a set of routes, one for each agent, covering all sub-tasks, as illustrated in Fig. 2d. We note that in future work, the GATR architecture could be easily extended, e.g. to other encoders tailored for min-max objectives (e.g. [dpn], which has separate embeddings for task partitioning and navigation), or decoders that output actions for agents in parallel to encourage collaborative behavior [parco].
@@ III-A Analysis for small problem sizes
We evaluate mission times against four benchmarks: the optimal and median population solutions (Opt-P and Med-P respectively), calculated via brute-force enumeration of the full population of possible solutions (e.g., 1,814,400 for 3 agents and 8 sub-tasks, from Equation 1, evaluation of which takes $\sim$1 minute for a single scenario on the GPU); the median of a sample of the population - 101 randomly-generated solutions (Med-S); and a near-optimal solution generated by the LKH3 solver [lkh3] (NOpt-LKH3). Opt-P and NOpt-LKH3 serve as the upper bounds on expected performance, whilst the two median benchmarks represent a lower bound as they simulate the expected performance of a randomly generated plan. We note that although it is feasible to obtain enumeration solutions for these small problem sizes, this is not possible for the larger problems we investigate in Section III-B, for which NOpt-LKH3 and Med-S are essential for evaluation instead.